01-31-07 Weekly meeting
24th Street 1211
01-31-07 Weekly meeting
3/2/2022 1:02:06 AM
3/1/2022 10:38:47 AM
There are no annotations on this page.
Document management portal powered by Laserfiche WebLink 9 © 1998-2015
All rights reserved.
Pages to print
Enter page numbers and/or page ranges separated by commas. For example, 1,3,5-12.
After downloading, print the document using a PDF reader (e.g. Adobe Reader).
Download electronic document
View plain text
Meeting Minutes <br /> Detailed,Grouped by Topic for each Meeting <br /> Item Meeting Item Description Resp Status Due Date Compl'd Cls'd <br /> 019-0 Reference 5/17/06-12/13/06 minutes for past Carl Tully(NBBJ); Yes <br /> 01 discussions. <br /> 1/3/07: Walkway design is completed. Should have <br /> civil details next week. <br /> 1/10/2007: CarIT is continuing to work on design <br /> walkway. He still has some questions to the Owner. <br /> 1/17/07: CarIT&MarcE to discuss Lighting and other <br /> minor issues after the meeting. New walkway design <br /> will have to be reviewed with regards to removing the <br /> engineering trailer at a later date. <br /> 1/24/07: Sparling will be looking into lighting <br /> requirements. One final design is determined. Will <br /> need to fall back and make sure Tom's trailer can be <br /> moved. <br /> 1/31/2007: Lydig received CP-35 and distributed to <br /> subs. Followup is required to make sure TomB can get <br /> his trailer out once walkway is installed. <br /> VCT Flooring <br /> 026-0 Reference 5/17/06-12/20/06 minutes for past Carl Tully(NBBJ) Yes <br /> 02 discussions. <br /> 1/3/07: CarIT and Diane reviewed in place application of <br /> tile at a well used commercial office space,and appears <br /> to be very durable. CarIT to send review and comments <br /> to MarcE for determination. <br /> 1/10/2007: CarIT needs to follow up with e-mail to <br /> MarcE on review. This is beginning to be a hot item;a <br /> decision needs to be put in place as soon as possible. <br /> This is becoming a hot item. There will need to be time <br /> for cost evaluation and this may be an issue with lead <br /> times for the flooring selected. <br /> 1/17/07: CarIT has to send correspondence to MarcE <br /> with recommendation and summary of decisions made. <br /> 1/24/07: CarIT forwarded memo with photos to MarcE <br /> to review with Barb and Vince. <br /> 1/31/07: Lydig received CP-36 and distributed to subs. <br /> Manufacturer pricing difference? CarIT to inquire and <br /> forward info to BobH to check sub pricing. <br /> OB Revisions <br /> 027-0 Reference 5/17/06-11/22/06 minutes for past Bob Heggenes No <br /> 02 discussions. (LCI) <br /> 12/7/2006: NBBJ sent BobH preliminary plans to <br /> provide ROM. <br /> 12/13/2006: BobH to contact Blythe and Ewing for <br /> feedback,and provide price and write up tomorrow for <br /> OB unit changes. <br /> 12/20/2006: All AHU's will be on site by the end of <br /> January-Phase I&II. I.H.would like to review impact of <br /> Imaging changes prior to kicking off any OB changes. <br /> 1/3/07: OB decision to remain on hold until Lydig <br /> Prolog Manager Printed on: 2/1/2007 Lydig_Projects Page 2 <br />
The URL can be used to link to this page
Your browser does not support the video tag.